# Estimating the linear relation between variables that are never jointly observed

### v3

@@ 7.1 Simulation design
We performed a series of simulations in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice. The number of animals observed per group is chosen for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for every group are also considered to be equal, i.e. $w_{1}=\cdots=w_{K}=\frac{1}{K}$. For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$. We simulate the predicted variable independently as $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}$, where $X^{\prime k}_{i}\sim X^{k}_{i}$ and $\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with regression parameters $\beta_{0}=1$ and $\beta_{1}=2$. It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}$ times. For each simulation, $N_{boot}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated. Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e. the proportion of intervals not including 0. Along with the bootstrap estimators, we also considered asymptotic confidence intervals in case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1. Additionally, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
@@
Figure 7.1: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
Throughout all simulations, we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired if the measurements are available only for a small number of animals, which is often the case for real experimental data. We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution. Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished one from another. We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e. $\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$. The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, hence the same level of $\sigma_{\epsilon}$ cannot be interpreted the same way for different values of $\sigma_{X}$. The variance of the noise can be expressed as follows: $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 7.1.
@@ 7.2 Results
The results of the simulation study are presented in Supplementary Figure C.1. Firstly, the results for the moment approach in the asymptotic and bootstrap cases are very similar, thus confirming the effectiveness of the bootstrap approach. Concerning the comparison between the bootstrap procedure and the naive approach, it can be observed that in general the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, at the expense of a slightly lower coverage rate. Whereas the empirical coverage rates in all cases are close to the nominal one (95%), the extent to which the amplitudes are smaller and the powers are bigger for the bootstrap estimators is very important in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance. This trend is further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches (the case with few groups and animals is presented in Figure 7.2). Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present. Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
@@
Figure 7.2: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$. The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X}$ and the noise to signal ratio $\rho$. The lines indicate the method used to estimate the confidence intervals: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
@@
Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X}$ and higher $\rho$. In most cases, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high dispersion, high noise level, few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes. On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals. This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
@@
Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
@@ 8 Application to real data
In order to illustrate the proposed estimators on real data, we studied the data previously mentioned in Section 1, obtained from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume. In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy. For one mouse cohort, the expressions of tho pro-inflammatory genes (IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a severity measure of radio-induced pulmonary lesions. In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung). A control condition is also considered, where the gene expressions are measured without prior radiation exposure. The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as a predictor and the septal thickening as as outcome. Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
Table 8.1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
@@
In order to ensure the comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation. The linear regression parameters were then estimated with three estimators the same way it was done in the simulation study in Section 7. The focus is placed on estimating the slope parameter $\beta_{1}$ in particular. The results are presented in Table LABEL:tab:vivo_results, containing the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance. On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (both genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result. The results are also consistent for all methods in the case without radiation exposure (control), where no significant relationship has been identified, as it is expected. On the other hand, we expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch. This is only the case for the bootstrap estimators, but not for the naive one. This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
Among the SBRT irradiation configurations, only the 3mm beam size has shown a significant correlation between inflammatory genes and septal thickening. These results are consistent with literature, indicating it as the beam size starting with which the long-term lesions appear (Bertho et al.,, 2020). Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend. These values are in general bigger in case of whole-thorax irradiaion and within the patch than for the ipsilateral or right lung for SBRT, which suggests a stronger correlation between the inflammatory process and lung injuries under high dose/volume irradiation conditions. These results, which are in line with the biological knowledge, could not have been established within the framework of classical statistical regression approaches due to non-simultaneous observations. This effect is illustrated in Figure 8.1 on the example of the linear model prediction made for the gene IL6.
Figure 8.1: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.

### v4

@@ 7.1 Simulation design
We performed a series of simulations to evaluate the finite sample performance of the proposed approaches on data resembling in vivo data from real experiments on mice. The number of animals observed per group is chosen for simplicity to be $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for each group are also assumed to be equal, i.e. $w_{1}=\cdots=w_{K}=\frac{1}{K}$. For each animal $i\in\{1,\dots,n\}$ and each subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$. We simulate the predicted variable independently as $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}$, where $X^{\prime k}_{i}\sim X^{k}_{i}$ and $\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with regression parameters $\beta_{0}=1$ and $\beta_{1}=2$. By making $X^{k}_{i}$ and $X^{\prime k}_{i}$ independent, we recreate the situation where the predictor and predicted variables are not observed simultaneously.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}$ times. For each simulation, $N_{boot}$ bootstrap samples of size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then the moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are computed. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are computed. Based on the $N_{boot}$ estimates, we compute 95% confidence intervals using the quantile function from the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, which we use to calculate the following quantities of interest: the coverage rate of the intervals, their average amplitude, and the power, i.e. the proportion of intervals that do not contain 0. In addition to the bootstrap estimators, we also considered asymptotic confidence intervals in the case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1. In addition, we estimate the confidence intervals for the parameter estimators of the regression on the means per group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the standard error of the estimator follows a Student’s t-distribution:
@@
Finally, we considered the case where the predictor and the predicted variable are observed simultaneously, i.e. $X^{k}_{i}$ and $Y^{k}_{i}$ are such that $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{k}_{i}+\epsilon^{k}_{i}$. In this case the parameters are estimated with the classical linear regression approach, and the confidence intervals are obtained with a Student’s t-distribution.
Throughout all simulations we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether inference is significantly affected when measurements are only available for a small number of animals, which is often the case in real experimental data. We consider the number of groups $K\in\{4,10\}$, where 4 is the number of groups often observed in real data, and 10 is a higher number that may produce sufficiently good results with the naive approach of approximating confidence intervals with the Student distribution. Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished from each other. We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to less overlap between groups and the second to more overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ which controls the variance of the response to the variance of the noise ratio, i.e. $\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$. The choice of adjusting the signal-to-noise ratio rather than the amount of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, so the same level of $\sigma_{\epsilon}$ cannot be interpreted in the same way for different values of $\sigma_{X}$. The variance of the noise can be expressed as follows $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated in Figure 3.
Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
@@ 7.2 Results
The results of the simulation study are shown in Supplementary Figure 7. Firstly, the results for the moment approach are very similar in the asymptotic and bootstrap cases, confirming the effectiveness of the bootstrap approach. When comparing the bootstrap procedure with the naive approach, it can be observed that the bootstrap estimators generally produce confidence intervals with smaller average amplitudes and higher power at the expense of a slightly lower coverage rate. Whereas the empirical coverage rates are close to the nominal one (95%) in all cases, the extent to which the average amplitudes are smaller and the powers are greater for the bootstrap estimators is very important in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance. This trend is further strengthened by the parameter encoding the number of groups: while the overall results worsen with a decrease in either the number of animals or the number of groups, it is the case with a small number of groups that shows the greatest difference between the approaches (the case with few groups and animals is shown in Figure 4). Indeed, in almost all cases within the tables with $K=4$ we observe that the average amplitudes are approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important because lower power implies a higher probability of not detecting a significant relationship between the predictor and the predicted variables, if it is indeed present. Overall, these results imply that the proposed bootstrap estimators are more effective when the experimental design involves a small number of groups.
@@
Figure 4: a) Coverage rates, b) average amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$. The columns of the tables indicate simulation scenarios with different combinations of parameters: scenario S1 with lower group overlap ($\sigma^{2}_{X}=0.75$) and higher signal-to-noise ratio ($\rho=1.1$), S2 with higher group overlap ($\sigma^{2}_{X}=2$) and higher signal-to-noise ratio ($\rho=1.1$), S3 with lower group overlap ($\sigma^{2}_{X}=0.75$) and lower signal-to-noise ratio ($\rho=1.01$), and S4 with higher group overlap ($\sigma^{2}_{X}=2$) and lower signal-to-noise ratio ($\rho=1.01$). The lines indicate the method used to estimate the confidence intervals: ”mm (asymp)” stands for the method of moments with asymptotic confidence intervals, ”mm (boot)” for the method of moments with bootstrap, ”ot (boot)” for the optimal transport method with bootstrap, ”mm (student)” for the naive linear regression on means approach based on Student’s distribution, and ”simultaneous” for the classical linear regression estimation in the case where the predictor and the predicted variable are observed simultaneously.
@@
Regarding the remaining two parameters, as expected, the best results are generally obtained with lower $\sigma^{2}_{X}$ and higher $\rho$. In most cases, the results for the naive and bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high overlap, high noise, few groups and few animals, where all estimators fail drastically: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in average amplitudes. On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are above 90% while those of the naive estimator are below 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but less overlap, in the second case a high level of overlap is compensated by a higher number of animals. This means that if the underlying distributions per group are characterized by a reasonable amount of overlap, or if a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
Figure 5: Distributions of amplitudes of confidence intervals obtained with different methods based on 500 simulations under scenarios S2 (a) and S3 (b), with $K=4$ and $n=10$.
Concerning the comparison with the simultaneous case, as expected, the latter globally produces better results than any method in the context of non-simultaneous design. In particular, average amplitudes are systematically 10%-40% smaller compared to our methods in the non-simultaneous case. However, the difference seems to be less important than that between our approaches and the naive approach. Boxplots of the amplitude distributions for the most challenging combination of $K$ and $n$ under the most interesting scenarios (S2 and S3) are shown in Figure 5. The figure confirms the fact that the amplitude distributions obtained with our methods are more similar to those obtained in the simultaneous case than to those obtained with the naive approach in the non-simultaneous case. Moreover, the powers are almost as high as those obtained in the simultaneous case in 3 out of 4 scenarios. The only scenario where this difference is striking is the one with little noise and high group overlap. It can be concluded that if the groups are separable, in the case where the predictor and the predicted variable are not observed simultaneously, with our approach we can detect a significant relationship with a success comparable to the case when the variables are observed simultaneously. It should be noted that if the data are too noisy, all methods fail to detect the significance in all cases.
@@
Finally, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller average amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in cases with higher group overlap $\sigma^{2}_{X}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator is likely to produce better results in terms of power than the method of moments estimator when the bias is corrected.
@@ 8 Application to real data
To illustrate the proposed estimators on real data, we examined the data mentioned in Section 1, obtained from experiments conducted in mice to assess the adverse effects induced in the context of different irradiated volumes. In these experiments (see Bertho et al., (2020) for a more detailed presentation), mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy to the left lung, or whole thorax irradiation (WTI) at 19 Gy. For one cohort of mice, the expression of pro-inflammatory genes (IL6 and TNF) was measured, and for the other cohort, the thickness of the alveolar septum was measured as a measure of the severity of radiation-induced lung lesions. In the case of SBRT, measurements were taken at multiple sites: the irradiated patch (within the irradiation field), the remaining part of the left lung, referred to as the ipsilateral lung, and the right lung (contralateral lung). A control condition is also included where gene expression is measured without prior exposure to radiation. The goal of this statistical analysis is to determine whether there is a statistical association between gene expression as a predictor and septal thickening as an outcome. Our approach is applied because the variables are measured on different animals, but within each irradiation condition there are common groups in terms of measurement time points.
To ensure comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation. The linear regression parameters were then estimated with three estimators in the same way as in the simulation study in Section 7. The focus is on the estimation of the slope parameter $\beta_{1}$. The results are presented in Table 1, which contains the estimates of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator and the corresponding test result for the significance of the estimated relationship.
Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
@@
Considerable differences can be observed between the results for the naive and the bootstrap estimators with respect to the significance found. On the one hand, the relationship between the pro-inflammatory genes and septal thickening was identified by all methods in the case of whole thorax irradiation (both genes for the bootstrap estimators and only one gene for the naive estimator), which is an expected result. The results are also consistent for all methods in the case of no radiation exposure (control), where no significant association was identified, as expected. On the other hand, we expect to identify a strong correlation in the case of measurements taken directly from the irradiated patch. This is the case only for the bootstrap estimators, but not for the naive one. This result is in accordance with the results obtained with simulated data: the confidence intervals are often overestimated with the naive approach, which can lead to false negatives in terms of significance.
Among the SBRT irradiation configurations, only the 3 mm beam size showed a significant correlation between inflammatory genes and septal thickening. These results are consistent with the literature, indicating that this is the beam size from which the long-term lesions appear (Bertho et al.,, 2020). Several significant associations were identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
Finally, in the cases where a significant relationship was detected, the estimated values of the slope are always positive, indicating a general radio-induced upregulation trend. These values are generally greater for whole thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, suggesting a stronger correlation between the inflammatory process and lung injury under high dose/volume irradiation conditions. These results, which are in line with biological knowledge, could not have been obtained using classical statistical regression approaches due to non-simultaneous observations. This effect is illustrated in Figure 6 using the example of linear model prediction for the gene IL6.
Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.